OpenStudy (anonymous):

Let \(\Large f(x) = log(1-x)\) a) Write down all higher derivatives \(f^{(k)}(x)\) for \(k\geq 1\) in closed form (i.e. as rational functions)

OpenStudy (anonymous):

\[\frac{1}{1-x}\] \[\frac{1}{(1-x)^2}\] \[\frac{2}{(1-x)^3}\] \[\frac{6}{(1-x)^4}\] \[...\]

OpenStudy (anonymous):

pattern is clear yes? power in the denominator increases by 1 each time the numerator may or may not be clear, but one more should suffice \[\frac{24}{(1-x)^5}=\frac{4\times 6}{(1-x)^5}=\frac{4\times 3\times 2}{(1-x)^5}\]

OpenStudy (anonymous):

@satellite73 thank you i was just asking about it, you read my mind :) just numerator i could not figure out .. how we get 1 2 6 i would say 1x2x3 but .. i dont know :)

OpenStudy (anonymous):

Now i get it :) thank you...